OpenStudy (anonymous):

12^-3 x 12^10 x 12^0? simplify

OpenStudy (ivettef365):

that is equal to 12^(-3+10+0) =

OpenStudy (dls):

\[\Huge (x)^a \times (x)^b \times (x)^c = (x)^{a+b+c}\]

OpenStudy (anonymous):

1/12^3 x 12^10 x 1= 12^7

OpenStudy (dls):

x=12 a=-3 b=10 c=0

OpenStudy (anonymous):

so the answer would be 12^7??